OpenStudy (anonymous):

solve the radical equation √7x+44=x

OpenStudy (anonymous):

x = 65.3955

OpenStudy (anonymous):

i need work..

OpenStudy (anonymous):

√7x+44=x 7x+44=x^2 x^2-7x-44=0 x^2-11x+4x-44=0 x(x-11)+4(x-11)=0 (x+4)(x-11)=0 x=-4,11 but for -4 the expression becomes complex so x=11 is the result